# Metaplectic time-frequency representations

### v1

@@ 2.1. The 20th century
A key effort to identify a common thread unifying time-frequency representations is due to Cohen [7], who observed how most of the above-mentioned time-frequency representations can be written as convolutions between the ambiguity function and a distribution, later named kernel by Claasen and Mecklenbräuker [6]. The family of such time-frequency representations is known as Cohen’s class at the present. It is straightforward to observe that the Cohen’s class can be defined in terms of the Wigner distribution, [6]. so that a time-frequency representation $Q$ belongs to the Cohen’s class if
@@ 2.3. Metaplectic Wigner distributions
The recent contributions by Cordero and Rodino yielded the definition of the $\mathcal{A}$-Wigner distributions in 2022 [17]. Their core observation was that many of the most important time-frequency representations defined in the last two centuries can be written letting a metaplectic operator on $L^{2}({\mathbb{R}^{2d}})$ acting on the tensors
@@ 3.2. Shift-invertiblity
We say that $W_{\mathcal{A}}$, or equivalently $\mathcal{A}$, is shift-invertible if $E_{\mathcal{A}}\in\operatorname*{GL}(d,\mathbb{R})$.
@@
    There exist $E\in\operatorname*{GL}(d,\mathbb{R})$, $C\in\operatorname*{Sym}(d,\mathbb{R})$, $\hat{\delta}\in\operatorname*{Mp}(d,\mathbb{R})$ and $c\in\mathbb{C}$, $|c|=1$, such that

### v2

@@ 2.1. The 20th century
A key effort to identify a common thread unifying time-frequency representations is due to Cohen [7], who observed how most of the above-mentioned time-frequency representations can be written as convolutions between the ambiguity function and a distribution, later named kernel by Claasen and Mecklenbräuker [6]. The family of such time-frequency representations is known as Cohen’s class at the present. It is straightforward to observe that the Cohen’s class can be defined in terms of the Wigner distribution, [6], so that a time-frequency representation $Q$ belongs to the Cohen’s class if
@@ 2.3. Metaplectic Wigner distributions
The recent contributions by Cordero and Rodino yielded the definition of the $\mathcal{A}$-Wigner distributions in 2022 [17]. Their core observation was that many of the most important time-frequency representations defined in the last two centuries can be written letting a metaplectic operator on $L^{2}({\mathbb{R}^{2d}})$ act on the tensors
@@ 3.2. Shift-invertiblity
We say that $W_{\mathcal{A}}$, or equivalently $\mathcal{A}$, is shift-invertible if $E_{\mathcal{A}}\in\operatorname*{GL}(2d,\mathbb{R})$.
@@
    There exist $E\in\operatorname*{GL}(2d,\mathbb{R})$, $C\in\operatorname*{Sym}(2d,\mathbb{R})$, $\hat{\delta}\in\operatorname*{Mp}(d,\mathbb{R})$ and $c\in\mathbb{C}$, $|c|=1$, such that